OpenStudy (kweenlexis):

Am i right???? Evaluate. 3−1⋅(4⋅6)⋅2−3 Enter your answer in the box. 0.1725??

OpenStudy (ageta):

3-1 * 24 * 2 -3 3-24*2-3 3-48-3 -48

OpenStudy (ageta):

Use PMDAS

OpenStudy (kweenlexis):

my bad i typed it wron its 3^-1 *(4.6)* 2^-3 @ageta

jhonyy9 (jhonyy9):

so 3^(-1) = ?

jhonyy9 (jhonyy9):

do you know it ?

jhonyy9 (jhonyy9):

3^(-1) = 1/3

OpenStudy (kweenlexis):

i thought it wass 0.3

jhonyy9 (jhonyy9):

4*6=24 and 2^(-3) = ?

OpenStudy (kweenlexis):

0.125?

jhonyy9 (jhonyy9):

yes but you need evaluate this your posted exercise

OpenStudy (kweenlexis):

ok

jhonyy9 (jhonyy9):

3^(-1) *(4*6) *2^(-3) = (1/3) *24 *1/2^3 = 24/(3*8) = ?

OpenStudy (kweenlexis):

64?

jhonyy9 (jhonyy9):

24 divide (3*8) = ?

OpenStudy (kweenlexis):

1?

jhonyy9 (jhonyy9):

yes sure this is equal 1

jhonyy9 (jhonyy9):

bc. you get 24/24 = 1

jhonyy9 (jhonyy9):

do you understand it now ?

OpenStudy (kweenlexis):

where you get fraction from though?

jhonyy9 (jhonyy9):

a^(-1) = 1/a this is propertie of exponents

OpenStudy (kweenlexis):

oh

jhonyy9 (jhonyy9):

you dont learned it again ?

jhonyy9 (jhonyy9):

2^(-3) = 1/(2^3) = 1/8

jhonyy9 (jhonyy9):

ok. ?

OpenStudy (kweenlexis):

yeah i guess

jhonyy9 (jhonyy9):

so was my pleasure good luck bye bye